Mrs. Mone wants to build a new rectangular playground for the kindergarten and use at most 350 feet of anchor fencing. He wants the length of the playground to be 50 feet longer than the three times the width. Find the dimensions of the playground to the nearest foot
Answer by ankor@dixie-net.com(22740)

Mr. Mone wants to build a new rectangular playground for the kindergarten and use at most 350 feet of anchor fencing.
He wants the length of the playground to be 50 feet longer than the three times the width.
Find the dimensions of the playground to the nearest foot
L = 3w + 50
:
2L + 2w = 350
simplify, divide by 2
L + w = 175
Replace L with (3w+50)
(3w+50) + w = 175
4w = 175 - 50
4w = 125
w = 125/4
w = 31.25 or 31 ft is the width to the nearest foot
Find L
L = 3(31) + 50
L = 93 + 50
L = 143 ft is the length
:
Find the actual amt of fence required with these dimensions
2(143) + 2(31) =
286 + 62 = 348 ft Which about the max to the nearest foot